Barton Industries expects that its target capital structure for raising funds in the future for its capital budget will consist of 40% debt, 5% preferred stock, and 55% common equity. Note that the firm's marginal tax rate is 40%. Assume that the firm's cost of debt, rd, is 6.6%, the firm's cost of preferred stock, rp, is 6.1% and the firm's cost of equity is 10.6% for old equity, rs, and 11.34% for new equity, re. What is the firm's weighted average cost of capital (WACC1) if it uses retained earnings as its source of common equity? Round your answer to 3 decimal places. Do not round intermediate calculations.
%

What is the firm’s weighted average cost of capital (WACC2) if it has to issue new common stock? Round your answer to 3 decimal places. Do not round intermediate calculations.
%
Answer #1

After-tax cost of debt=6.6*(1-tax rate)

=6.6*(1-0.4)=3.96%

WACC=Respective costs*Respective weight

a.WACC=(3.96*0.4)+(6.1*0.05)+(10.6*0.55)

=7.719%

b.WACC=(3.96*0.4)+(6.1*0.05)+(11.34*0.55)

=8.126%